OpenStudy (adamaero):

Integral from -pi/2 to pi/2 ( 16 + 24sinx + 9sin^2(x) ) ? always MEDAL

OpenStudy (adamaero):

In the solution manual I don't get why it cities about odd and even functions then magically, the middle term (24sinx) goes away with the problem exactly the same.

OpenStudy (adamaero):

\[1/2\int\limits_{-\pi/2}^{\pi/2}16 +24sinx + 9\sin ^{2}x dx\]

OpenStudy (adamaero):

I'm in the area of polar coordinates chapter.

OpenStudy (adamaero):

Do not solve the integral! I'm just asking how the 24sinx went away.

OpenStudy (adamaero):

@SolomonZelman

OpenStudy (adamaero):

@wio

OpenStudy (jhannybean):

The only way I could think of it going away is integrating the portion \(\int_{\pi/2}^{\pi/2} 24\sin(x)\)

OpenStudy (adamaero):

ya, actually I think they just didn't show that step

OpenStudy (adamaero):

should've just continued and solved it

OpenStudy (adamaero):

You'd be interested in this: http://openstudy.com/study#/updates/548912eae4b0209d5c4e0cbc

OpenStudy (solomonzelman):

16 + 24sinx + 9sin^2(x) = (4+3sin(x))(4+sin(x))

OpenStudy (solomonzelman):

go ahead, lol

OpenStudy (jhannybean):

\[\large \int_{-\pi/2}^{\pi/2}16dx+\int_{-\pi/2}^{\pi/2} 24\sin(x)dx+\int_{-\pi/2}^{\pi/2}9\sin^2(x)\]

OpenStudy (solomonzelman):

Jhanny, it is a perfect square trinomial.

OpenStudy (jhannybean):

Ooh it is!! I see that now.

OpenStudy (solomonzelman):

\[\int\limits_{-\pi /2}^{\pi/2} ( 3 \sin x +4 ) ^2~dx\]

OpenStudy (solomonzelman):

Actually, that doesn't necessarily simplify in this case.

OpenStudy (jhannybean):

Yeah it's not a simple enough function lol

OpenStudy (jhannybean):

You'd need a u sub and stuff :(

OpenStudy (jhannybean):

Maybe.

OpenStudy (adamaero):

no, half angle identity

OpenStudy (solomonzelman):

I don't know why the mayde it such a perfect square trinomial though. Im have not so much time right now, I wish i did....

OpenStudy (solomonzelman):

cosine or sine, why does that make a difference, perhapsm doing it just term by term as you did is the best way.

OpenStudy (solomonzelman):

I got to go, sorry guys.

OpenStudy (jhannybean):

\[\sin(2x) = \frac{1}{2} +\frac{1}{2}(1-\cos(2x))\]